Evulate ∫ (2x^2 + 3y^2)sc, where C is the curve
C
x(t)= at^2, y(t) = 2at, 0≤t≤1
1
Expert's answer
2021-07-19T05:51:53-0400

"∫ (2x^2 dx+ 3y^2dy)\\
x(t)= at^2 \implies dx=2atdt\\
y(t) = 2at\implies dy =2adt\\
∫ (2x^2 dx+ 3y^2dy)=∫_0^12(at^2)^2*2atdt+3(2at)^22adt\\
∫_0^14a^3t^5dt+∫_0^112a^3t^2dt\\
∫_0^14a^3t^5dt+∫_0^112a^3t^2dt\\
\frac{2}{3}a^3+4a^3\\
\frac{14a^3}{3}"